# Promoted current-induced spin polarization in inversion symmetry broken topological insulator thin films

### v1

@@ V Vertex corrections
By substituting Eqs. (22) into Eq. (18), the velocity correction $\delta$ can be computed numerically.
@@
VI Conclusions
We have developed a comprehensive theoretical framework to investigate current-induced spin polarization in topological insulator thin films with broken inversion symmetry. Using the Kubo formalism within the self-consistent Born approximation and including vertex corrections, we analyzed how the transverse spin susceptibility $\chi_{yx}$ is influenced by key factors, such as chemical potential ($\mu$), potential difference ($V$), hybridization parameters ($\Delta_{0}$ and $\Delta_{1}$), and disorder.
@@ VI Conclusions
Overall, our findings provide deeper insights into the current-induced spin polarization in TI thin films. They also highlight the interplay between chemical potential, hybridization, and external electric fields in controlling spin polarization in TI thin films, thereby opening new possibilities for advanced spintronic applications such as spin transistors and nonvolatile memory devices.

### v2

@@ 5 Vertex corrections
By substituting Eqs. (22) into Eq. (19), the velocity correction $\delta$ can be computed numerically (see the Appendix).
@@
6 Conclusions
We have developed a comprehensive theoretical framework to investigate CISP in topological insulator thin films with broken inversion symmetry. Using the Kubo formalism within the self-consistent Born approximation and including vertex corrections, we analyzed how the transverse spin susceptibility $\chi_{yx}$ is influenced by key factors, such as chemical potential ($\mu$), potential difference ($V$), hybridization parameters ($\Delta_{0}$ and $\Delta_{1}$), and disorder.
@@ 6 Conclusions
Although our study is based on a generic model, for the relevant scales of the parameters, we refer to recent first-principles modeling of Bi2Se3, Bi2Te3, and Sb2Te3 thin films [100, 93]. For Bi2Se3 thin films with thicknesses 3–6 QLs, the fitted momentum-independent gap takes values $\Delta_{0}\approx 3$120120 meV (corresponding to $\Delta_{0}\approx 0.006$$0.10$ in dimensionless units), while the momentum-dependent gaps are $\Delta_{1}\approx-16$ eVÅ2 (corresponding to $\Delta_{1}\approx-2.0$). For Bi2Te3, the hybridization gap is smaller, $\Delta_{0}\approx 1.5$1414 meV ($\Delta_{0}\approx 0.005$$0.07$), with $\Delta_{1}\approx-28$ to 30-30 eVÅ2 ($\Delta_{1}\approx-5$ to 7-7). Sb2Te3 films show somewhat larger gaps at small thickness, $\Delta_{0}\approx 6$3232 meV ($\Delta_{0}\approx 0.005$$0.05$), and $\Delta_{1}\approx-13$ to 16-16 eVÅ2 ($\Delta_{1}\approx-1.0$ to 1.3-1.3). Remarkably, for ultrathin films with a thickness of 2 QLs, in the cases of Bi2Se3 and Sb2Te3, the above-mentioned values can be up to one order of magnitude larger [100]. Thus, the parameter windows we employ map directly onto the experimentally accessible regimes.
Finally, our predicted features of current-induced spin polarization can be directly probed by current experimental statues such as spin potentiometry [101] and spin-torque ferromagnetic resonance [102]. These methods have already been applied to Bi2Se3 and Bi2Te3 films. Circularly polarized photoconductivity [75] and nonlocal spin–charge conversion setups [69] provide additional approaches to probe the role of inversion asymmetry and thickness.
@@ Acknowledgement
Appendix: Derivation of the velocity vertex corrections
In the ladder approximation, the dressed velocity $\tilde{v}_{x}$ is determined by
$\tilde{v}_{x}=v_{x}+n_{i}V_{0}^{2}\int\!\frac{d^{2}k}{(2\pi)^{2}}\,G^{+}\,\tilde{v}_{x}\,G^{-}.$ (23)
We expand the dressed vertex in the 16-matrix basis $\{\tau_{i}\!\otimes\!\sigma_{j}\}$. The in-plane rotational symmetry, time-reversal symmetry, and scalar disorder restrict the allowed channels. For the present case only two basis elements mix under the ladder kernel:
$\tilde{v}_{x}=a\,(\tau_{z}\otimes\sigma_{y})+b\,(\tau_{z}\otimes\sigma_{x})-\frac{2k_{x}\Delta_{1}}{\hbar}\,(\tau_{x}\otimes\sigma_{0}).$ (24)
Projecting Eq. (23) onto these two channels leads to the $2\times 2$ linear system
$\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}v\\ 0\end{pmatrix}+\begin{pmatrix}\mathcal{I}_{yy}&\mathcal{I}_{yx}\\ \mathcal{I}_{xy}&\mathcal{I}_{xx}\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}$, (25)
where the kernels $\mathcal{I}_{\alpha\beta}$ are defined by
$\mathcal{I}_{\alpha\beta}=n_{i}V_{0}^{2}\int\frac{d^{2}k}{(2\pi)^{2}}\,\mathrm{Tr}\!\left[\mathcal{P}_{\alpha}\,G^{+}\,\mathcal{P}_{\beta}\,G^{-}\right].$ (26)
Here $\mathcal{P}_{\alpha}$ are projection operators onto the relevant basis elements:
$\mathcal{P}_{yy}$ $=\tfrac{1}{4}(\tau_{z}\otimes\sigma_{y}),\qquad\mathcal{P}_{xx}=\tfrac{1}{4}(\tau_{z}\otimes\sigma_{x}),$ (27)
$\mathcal{P}_{yx}$ $=\tfrac{1}{4}(\tau_{z}\otimes\sigma_{x}),\qquad\mathcal{P}_{xy}=\tfrac{1}{4}(\tau_{z}\otimes\sigma_{y}).$ (28)
Because the system is rotationally invariant in the xxyy plane, the integrals reduce to
$\mathcal{I}_{\alpha\beta}=\frac{n_{i}V_{0}^{2}}{2\pi}\int_{0}^{\infty}\!\frac{k\,dk}{(2\pi)}\,\langle\mathrm{Tr}[\mathcal{P}_{\alpha}G^{+}(\mu,k,\theta)\,\mathcal{P}_{\beta}G^{-}(\mu,k,\theta)]\rangle_{\theta},$ (29)
where $\langle\cdots\rangle_{\theta}$ denotes averaging over the polar angle $\theta$ of $\mathbf{k}$.
Using the spectral representation of the Green’s function,
$G^{\pm}=\sum_{n=1}^{4}\frac{|u_{n,\mathbf{k}}\rangle\langle u_{n,\mathbf{k}}|}{\mu-E_{n}(k)\pm i\eta},$ (30)
with eigenvalues
$E_{l,p}(k)=l\sqrt{\Delta(k)^{2}+(vk+pV)^{2}},\quad l,p=\pm 1,$ (31)
and eigenvectors $|u_{n,\mathbf{k}}\rangle$ the kernels become
$\mathcal{I}_{\alpha\beta}=\frac{n_{i}V_{0}^{2}}{(2\pi)^{2}}\sum_{n,m}\int_{0}^{\infty}\!k\,dk\,\frac{\langle u_{n,\mathbf{k}}|\mathcal{P}_{\alpha}|u_{m,\mathbf{k}}\rangle\,\langle u_{m,\mathbf{k}}|\mathcal{P}_{\beta}|u_{n,\mathbf{k}}\rangle}{(\mu-E_{n}(k)+i\eta)(\mu-E_{m}(k)-i\eta)}.$ (32)
For numerical purposes, the angular average reduces most cross terms to zero by symmetry, leaving only diagonal contributions in $\sigma_{y}$ and $\sigma_{x}$ channels. The $k$-integrals are convergent and can be evaluated numerically. In practice, $\mathcal{I}_{xy}\simeq\mathcal{I}_{yx}\simeq 0$, so the system (25) reduces approximately to $a=v_{f}/(1-\mathcal{I}_{yy})$ and $b\simeq 0$, with $a=v_{f}/(1+\delta)$.